OpenStudy (anonymous):

Limit as x->0 (tan2x/x)

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{\tan2x}{x}=\lim_{x\to0}\frac{\frac{\sin2x}{\cos2x}}{x}\cdot\frac{2}{2}=2\left(\lim_{x\to0}\frac{1}{\cos x}\right)\left(\lim_{x\to0}\frac{\sin2x}{2x}\right)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome!